Differentiate with respect to the variable


y= sin3t, find dy/dt


v= 11sin4t - 5cos5t, find dv/dt


1
Expert's answer
2021-04-15T06:56:57-0400

Solution. Using the rules of differentiation and the table of derivates we get


"\frac{dy}{dt}=3cos(3t)"

"\frac{dv}{dt}=11\times4\times cos(4t)-5\times 5 \times (-sin(5t))=44cos(4t)+25sin(5t)"

Answer.


"\frac{dy}{dt}=3cos(3t)"


"\frac{dv}{dt}=44cos(4t)+25sin(5t)"